Westerville Company reported the following results from last year’s operations: Sales $ 1,200,000 Variable expenses 320,000 Contribution margin 880,000 Fixed expenses 640,000 Net operating income $ 240,000 Average operating assets $ 600,000 At the beginning of this year, the company has a $150,000 investment opportunity with the following cost and revenue characteristics: Sales $ 240,000 Contribution margin ratio 50 % of sales Fixed expenses $ 84,000 The company’s minimum required rate of return is 15%.

10-a. If Westerville’s chief executive officer will earn a bonus only if her ROI from this year exceeds her ROI from last year, would she pursue the investment opportunity? Yes No

10-b. Would the owners of the company want her to pursue the investment opportunity? Yes No

11. What is last year’s residual income?

12. What is the residual income of this year’s investment opportunity?

13. If the company pursues the investment opportunity and otherwise performs the same as last year, what residual income will it earn this year?
Answer #1

Below question based on multiple formula . I used all Alphabetic to derived FORMULA ( Like a, b c .... etc..)  

Residual Income last year , current year plus at consolidated level = formula based calculation - Updated in below answer sheet .

Any understanding issue , let me know .

Westerville Compaby reported
Amount ($)
Sales (a)       12,00,000
Variable cost(b)          3,20,000
Contribution Margin (a-b)=c          8,80,000
Fixed Expenses (d)          6,40,000
Net Operating Income (c-d)=e          2,40,000
Last year Margin====(e/a) 20%
Average Operating cost-f          6,00,000
Last year turnover (a/f)                  2.00 Times
ROI (e/f) 40%
Initial Investment(J)          1,50,000
Sales (I)          2,40,000
Contribution Margin ( 50% of Sales)(g)          1,20,000
Fixed Expenses (h)             84,000
Margin(g-h)== K             36,000
Margin %(k/i) 15%
Turnover (I/J)                  1.60 Times
ROI(K/J) 24%
Company Minimum Required return 15%
Total Sales (a+i)= L       14,40,000
Total Margin(e+k)=M          2,76,000
Margin %(M/L) 19%
Operating Asset ( f+j)= N          7,50,000
Turnover (L/N)                  1.92 Times
ROI(M/N) 37%
Last Year residual Income 150000
Margin - Minimum Return on Asset
('240000-15%*600000
Residual income current Year 13500
Margin - Minimum Return on Asset
('36000-15%*1,50,000
Consolidated level Residual Income 163500
Margin - Minimum Return on Asset
('276000-15%*750000)